# Stochastic gradient descent estimation of generalized matrix factorization models with application to single-cell RNA sequencing data

### v1

@@ 2 Model specification
The multivariate generalized linear model in (3) is non-identifiable being invariant with respect to rotation, scaling, and sign-flip transformations of $\mathbf{U}$ and $\mathbf{V}$. Then, to enforce the uniqueness of the matrix decomposition, we need to impose additional identifiability constraints. Some of the most common choices in the literature involve the following equivalent parameterizations:
@@
    Orthonormal scores: $\mathbf{U}$ has orthonormal columns, $\mathbf{U}^{\top}\mathbf{U}=\mathbf{I}_{d}$, $\mathbf{V}$ has orthogonal columns, $\mathbf{V}^{\top}\mathbf{V}=\boldsymbol{\Delta}^{2}$, the first non-zero element of each column of $\mathbf{V}$ is positive;
  • (B)
    Orthonormal loadings: $\mathbf{U}$ has orthogonal columns, $\mathbf{U}^{\top}\mathbf{U}=\boldsymbol{\Delta}^{2}$, $\mathbf{V}$ has orthonormal columns, $\mathbf{V}^{\top}\mathbf{V}=\mathbf{I}_{d}$, the first non-zero element of each column of $\mathbf{V}$ is positive;
  • (C)
    Standardizes scores: $\mathbf{U}$ has standardized columns, say $\mathbf{1}_{n}^{\top}\mathbf{U}/n=\mathbf{0}_{d}$ and $\mathbf{U}^{\top}\mathbf{U}/n=\mathbf{I}_{d}$, and $\mathbf{V}$ is lower triangular with positive diagonal entries.
Here, $\boldsymbol{\Delta}^{2}$ denotes the diagonal matrix collecting all the non-zero singular values of $\mathbf{U}\mathbf{V}^{\top}$ in decreasing order, $\mathbf{I}_{d}$ denotes the $d\times d$ identity matrix, $\mathbf{0}_{d}$ is the $d\times 1$ vector of zeros, and $\mathbf{1}_{n}$ is the $n\times 1$ vector of ones. Notice that any unconstrained estimate of $\mathbf{U}$ and $\mathbf{V}$ can be easily projected into the constrained space identified by (A), (B), or (C) by post-processing; see, e.g., Kidziński et al., (2022) and Wang and Carvalho, (2023). The choice of the parametrization typically depends on the specific application and the desired interpretation of the score and loading matrices. Unless explicitly stated otherwise, we will assume that parametrization (A) is used throughout.
Following the naming convention introduced by Kidziński et al., (2022), we refer to the model specification presented so far in (1)–(3) as generalized matrix factorization (GMF). Alternative nomenclatures, such as exponential family principal component analysis (EPCA, Collins et al.,, 2001; Mohamed et al.,, 2008; Li and Tao,, 2010), generalized low-rank models (GLRM, Udell et al.,, 2016), generalized linear latent variable model (GLLVM, Niku et al.,, 2017; Hui et al.,, 2017), generalized principal component analysis (glmPCA, Townes et al.,, 2019), deviance matrix factorization (DMF, Wang and Carvalho,, 2023), or generalized factor model (GFM, Liu et al.,, 2023), can also be found in the literature. Straightforward extensions of the GMF specification include pseudo-likelihood models, with $f(y_{ij};\boldsymbol{\psi})=\exp\{-L_{\phi}(y_{ij},\eta_{ij})\}$ being the negative exponential of a loss function $L_{\phi}(\cdot,\cdot):\mathcal{Y}\times\mathbb{R}\rightarrow\mathbb{R}_{+}$. Also, vector generalized estimating equations for overdispersed data are a particular case of this more general setup.
@@ 2.1 Penalized likelihood estimation
$\sum_{i=1}^{n}\log\int_{\mathbb{R}^{d}}\bigg{[}\prod_{j=1}^{m}f(y_{ij}|\mathbf{u}_{i:})\bigg{]}f(\mathbf{u}_{i:})\,\textrm{{d}}\mathbf{u}_{i:},$ (4)
@@
with $\Psi$ being the parameter space of $\boldsymbol{\psi}$, which incorporate the identifiability constrains in (A1)–(A2), and $\ell_{\lambda}(\boldsymbol{\psi};\mathbf{y})$ denoting the penalized negative log-likelihood function. The latter is given by
@@
where $\lambda>0$ is a regularization parameter and $\|\cdot\|_{F}^{2}$ denotes the Frobenious norm.
@@ 2.2 Related models
Then, in the GMF notation, PCA can be obtained by assuming a Gaussian distribution for the data matrix, an identity link function for the mean and no regression effects in the linear predictor, namely $(\mathbf{X}\mathbf{B}^{\top})_{ij}=0$ and $(\boldsymbol{\Gamma}\mathbf{Z}^{\top})_{ij}=0$. Generalizations of PCA, such as Binary PCA (Schein et al.,, 2003; Lee et al.,, 2010; Landgraf,, 2015; Song et al.,, 2019) and Poisson PCA (Kenney et al.,, 2021; Virta and Artemiou,, 2023), are included in the GMF framework.
@@ 3 Estimation algorithm
For the sake of exposition, throughout this section, we assume without loss of generality that $\mathbf{B}=\mathbf{0}$ and $\boldsymbol{\Gamma}=\mathbf{0}$ or, equivalently, that $\mathbf{U}_{\star}$ and $\mathbf{V}_{\star}$ are partially known matrices incorporating the regression effects and the covariate matrices. In the following, we suppress the $\star$ subscript to lighten the notation. Moreover, we consider $\phi$ as a known fixed parameter, noting that, if unknown, it can be estimated iteration–by–iteration using either a method of moments or via maximum likelihood. More details are provided in the Appendix. Finally, we introduce the matrices $\accentset{\mbox{\large.}}{\mathbf{D}}=\{\accentset{\mbox{\large.}}{D}_{ij}\}=\{\partial\ell_{\lambda}/\partial\eta_{ij}\}$ and $\accentset{\mbox{\large..}}{\mathbf{D}}=\{\accentset{\mbox{\large..}}{D}_{ij}\}=\{\partial^{2}\ell_{\lambda}/\partial\eta_{ij}^{2}\}$ for the derivatives of the deviance with respect to the linear predictor, where
$\accentset{\mbox{\large.}}{D}(y,\mu)=\frac{\partial\ell_{\lambda}}{\partial\eta}=\frac{w(y-\mu)}{\phi\,\nu(\mu)\,\accentset{\mbox{\large.}}{g}(\mu)},\qquad\accentset{\mbox{\large..}}{D}(y,\mu)=\frac{\partial^{2}\ell_{\lambda}}{\partial\eta^{2}}=\frac{w\,\alpha(\mu)}{\phi\,\nu(\mu)\{\accentset{\mbox{\large.}}{g}(\mu)\}^{2}},$
and $\alpha(\mu)=1+(y-\mu)\{\accentset{\mbox{\large.}}{\nu}(\mu)/\nu(\mu)+\accentset{\mbox{\large..}}{g}(\mu)/\accentset{\mbox{\large.}}{g}(\mu)\}$. Accordingly, we define $\accentset{\mbox{\large.}}{g}=\textrm{{d}}{g}/\textrm{{d}}\mu$, $\accentset{\mbox{\large..}}{g}=\textrm{{d}}^{2}{g}/\textrm{{d}}\mu^{2}$ and $\accentset{\mbox{\large.}}{\nu}=\textrm{{d}}\nu/\textrm{{d}}\mu$. Notice that the expected second-order derivative, i.e., the Fisher weight $\mathbb{E}(\accentset{\mbox{\large..}}{D}_{ij})$, just corresponds to the observed second-order differential $\accentset{\mbox{\large..}}{D}_{ij}$ with $\alpha(\mu_{ij})=1$, which is positive for any $y_{ij}$ and $\mu_{ij}$.
@@
$\frac{\partial^{2}\ell_{\lambda}}{\partial\mathbf{U}^{2}}=\accentset{\mbox{\large..}}{\mathbf{D}}\,(\mathbf{V}*\mathbf{V})+\boldsymbol{\Lambda}>0,\qquad\frac{\partial^{2}\ell_{\lambda}}{\partial\mathbf{V}^{2}}=\accentset{\mbox{\large..}}{\mathbf{D}}^{\top}(\mathbf{U}*\mathbf{U})+\boldsymbol{\Lambda}>0,$ (8)
where * is the Adamard product and $\boldsymbol{\Lambda}$ is a matrix of appropriate dimensions filled by $\lambda$.
Conditionally on $\mathbf{V}$, the left matrix equation in (7) can be decomposed into $n$ multivariate equations row–by–row, $\partial\ell_{\lambda}/\partial\mathbf{u}_{i:}=0$ ($i=1,\dots,n$), that can be solved independently in parallel. The existence and uniqueness of the solution of each row–equation are guaranteed under mild regularity conditions on the exponential family and the link function. In the same way, the right matrix equation in (7) can be split into $m$ independent vector equations, $\partial\ell_{\lambda}/\partial\mathbf{v}_{j:}=0$ ($j=1,\dots,m$), to be solved in parallel. See, e.g., Kidziński et al., (2022) and Wang and Carvalho, (2023) for a detailed discussion and derivation on (7) and (8). Notice that, in case of partially known matrices, say $\mathbf{U}_{\star}$ and $\mathbf{V}_{\star}$, the derivatives in (7) must be replaced by $\partial\ell_{\lambda}/\partial[\mathbf{A},\,\mathbf{U}]$ and $\partial\ell_{\lambda}/\partial[\mathbf{B},\,\mathbf{V}]$ to exclude from the differentiation the covariate matrices $\mathbf{X}$ and $\mathbf{Z}$.
@@ 3.1 Fisher scoring and quasi-Newton algorithms
The first, and most popular, algorithm introduced in the literature for finding the solution of (7) is the alternated iterative re-weighted least squares (AIRWLS, Collins et al.,, 2001; Li and Tao,, 2010; Risso et al.,, 2018; Kidziński et al.,, 2022; Wang and Carvalho,, 2023; Liu et al.,, 2023) method. It cycles between the conditional updates of $\mathbf{U}$ and $\mathbf{V}$ solving the equations in (7) in a row-wise manner, using standard Fisher scoring for GLMs (McCullagh and Nelder,, 1989). The resulting routine is statistically motivated, easy to implement and allows for efficient parallel computing.
@@
where $\{\rho_{t}\}$ is a sequence of learning rate parameters, $\boldsymbol{\Delta}_{\psi}^{t}$ is the search direction, while $\mathbf{G}_{\psi}^{t}=\partial\ell_{\lambda}^{t}/\partial\boldsymbol{\psi}$ and $\mathbf{H}_{\psi}^{t}=\partial^{2}\ell_{\lambda}^{t}/\partial\boldsymbol{\psi}^{2}$ denote respectively the first two derivatives of $\ell_{\lambda}(\boldsymbol{\psi};\mathbf{y})$ with respect to $\boldsymbol{\psi}$ evaluated at $\boldsymbol{\psi}^{t}$. Throughout, $\leftarrow$ stands for the assignment operator and the division is performed element-wise.
Exploiting the block structure of $\boldsymbol{\psi}$, the joint update (9) can be written in the coordinate-wise form
@@
where $\mathbf{G}_{\textsc{u}}$, $\mathbf{G}_{\textsc{v}}$, $\mathbf{H}_{\textsc{u}}$ and $\mathbf{H}_{\textsc{v}}$ can be obtained as in (7) and (8). Algorithm 1 provides a step-by-step pseudo-code description of the quasi-Newton method of Kidziński et al., (2022). To the best of our knowledge, this is the most efficient algorithm in the literature for the estimation of GMF models.
while convergence is not reached do
       Compute the likelihood derivatives
       $\boldsymbol{\eta}^{t}\leftarrow\mathbf{U}^{t}\mathbf{V}^{t\top}$;  $\boldsymbol{\mu}^{t}\leftarrow g^{-1}(\boldsymbol{\eta}^{t})$; $O(nmd)$
      $\accentset{\mbox{\large.}}{\mathbf{D}}^{t}\leftarrow\mathbf{W}*(\mathbf{Y}-\boldsymbol{\mu}^{t})/\big{\{}\phi^{t}\,\nu(\boldsymbol{\mu}^{t})*\accentset{\mbox{\large.}}{g}(\boldsymbol{\mu}^{t})\big{\}}$; $O(nm)$
       $\accentset{\mbox{\large..}}{\mathbf{D}}^{t}\leftarrow\mathbf{W}/\big{\{}\phi^{t}\,\nu(\boldsymbol{\mu}^{t})*\accentset{\mbox{\large.}}{g}(\boldsymbol{\mu}^{t})^{2}\big{\}}$; $O(nm)$
      Compute the batch gradients and update $\mathbf{V}$
       $\mathbf{G}_{\textsc{v}}^{t}\leftarrow\accentset{\mbox{\large.}}{\mathbf{D}}^{t\top}\mathbf{U}^{t}-\lambda\mathbf{V}^{t}$;  $\mathbf{H}_{\textsc{v}}^{t}\leftarrow\accentset{\mbox{\large..}}{\mathbf{D}}^{t\top}(\mathbf{U}^{t}*\mathbf{U}^{t})-\boldsymbol{\Lambda}$; $O(nmd)$
      $\boldsymbol{\Delta}_{\textsc{v}}^{t}\leftarrow-(\mathbf{G}_{\textsc{v}}^{t}/\mathbf{H}_{\textsc{v}}^{t})$;  $\mathbf{V}^{t+1}\leftarrow\mathbf{V}^{t}+\rho_{t}\,\boldsymbol{\Delta}_{\textsc{v}}^{t}$; $O(md)$
      Compute the batch gradients and update $\mathbf{U}$
       $\mathbf{G}_{\textsc{u}}^{t}\leftarrow\accentset{\mbox{\large.}}{\mathbf{D}}^{t}\,\mathbf{V}^{t}-\lambda\mathbf{U}^{t}$;  $\mathbf{H}_{\textsc{u}}^{t}\leftarrow\accentset{\mbox{\large..}}{\mathbf{D}}^{t}\,(\mathbf{V}^{t}*\mathbf{V}^{t})-\boldsymbol{\Lambda}$; $O(nmd)$
      $\boldsymbol{\Delta}_{\textsc{u}}^{t}\leftarrow-(\mathbf{G}_{\textsc{u}}^{t}/\mathbf{H}_{\textsc{u}}^{t})$;  $\mathbf{U}^{t+1}\leftarrow\mathbf{U}^{t}+\rho_{t}\,\boldsymbol{\Delta}_{\textsc{u}}^{t}$; $O(nd)$
Orthogonalize $\hat{\mathbf{U}}$ and $\hat{\mathbf{V}}$;
Algorithm 1 Pseudo-code description of the quasi-Newton algorithm of Kidziński et al., (2022). On the right, we report the computation complexity of each step.

### v2

@@ 2.1 Parameter identifiability
The multivariate generalized linear model in (3) is non-identifiable, allowing for column-space overlapping, and being invariant with respect to rotation, scaling, and sign-flip transformations of $\mathbf{U}$ and $\mathbf{V}$. Then, to enforce the uniqueness of the matrix decomposition, we need to impose additional identifiability constraints. First, we need to impose the orthogonality of the parameters with respect to the covariate column space:
@@
    $\mathbf{X}$ and $\mathbf{Z}$ are full-column rank matrices, moreover $\mathbf{X}^{\top}\boldsymbol{\Gamma}=\mathbf{0}$, $\mathbf{X}^{\top}\mathbf{U}=\mathbf{0}$, and $\mathbf{Z}^{\top}\mathbf{V}=\mathbf{0}$.
The first conditions ensure that $\mathbf{X}^{\top}\mathbf{X}$ and $\mathbf{Z}^{\top}\mathbf{Z}$ are non-singular, which is a standard requirement for regression models. Conditions $\mathbf{X}^{\top}\boldsymbol{\Gamma}=\mathbf{0}$ and $\mathbf{X}^{\top}\mathbf{U}=\mathbf{0}$ prevent $\boldsymbol{\Gamma}$ and $\mathbf{U}$ from spanning the same column space of $\mathbf{X}$, and similarly $\mathbf{Z}^{\top}\mathbf{V}=\mathbf{0}$ prevents $\mathbf{V}$ from spanning the same column space of $\mathbf{Z}$ (see, e.g., Liu and Zhong,, 2024).
Then, we must ensure the identifiability of $\mathbf{U}$ and $\mathbf{V}$ with respect to rotation, scaling, and sign-flip. To this end, some of the most common choices in the literature involve the following equivalent parameterizations:
  • (B1)
    $\mathbf{U}$ has orthogonal columns, $\mathbf{U}^{\top}\mathbf{U}=\boldsymbol{\Sigma}$, $\mathbf{V}$ has orthonormal columns, $\mathbf{V}^{\top}\mathbf{V}=\mathbf{I}_{d}$, the first non-zero element of each column of $\mathbf{V}$ is positive;
  • (B2)
    $\mathbf{U}$ has orthonormal columns, $\mathbf{U}^{\top}\mathbf{U}=\mathbf{I}_{d}$, $\mathbf{V}$ has orthogonal columns, $\mathbf{V}^{\top}\mathbf{V}=\boldsymbol{\Sigma}$, the first non-zero element of each column of $\mathbf{U}$ is positive;
  • (B3)
    $\mathbf{U}$ has standardized columns, say $\mathbb{V}\textrm{{ar}}(\mathbf{U})=\frac{1}{n}\mathbf{U}^{\top}(\mathbf{I}_{n}-\frac{1}{n}\mathbf{1}_{n}\mathbf{1}_{n}^{\top})\mathbf{U}=\mathbf{I}_{d}$, and $\mathbf{V}$ is lower triangular with positive diagonal entries.
Here, $\boldsymbol{\Sigma}=\textrm{{diag}}(\sigma_{1},\dots,\sigma_{d})$ denotes the diagonal matrix collecting all the non-zero singular values of $\mathbf{U}\mathbf{V}^{\top}$ in decreasing order, $\mathbf{I}_{d}$ denotes the $d\times d$ identity matrix, $\mathbf{0}_{d}$ is the $d\times 1$ vector of zeros, and $\mathbf{1}_{n}$ is the $n\times 1$ vector of ones. In Appendix A in the Supplementary Material, we prove that constraints (A), together with one of (B1), (B2), or (B3), are sufficient to ensure the identifiability of the model parameters. Notice that any unconstrained estimate of $\mathbf{B}$, $\boldsymbol{\Gamma}$, $\mathbf{U}$ and $\mathbf{V}$ can be easily projected into the constrained space induced by the identifiability restrictions via post-processing; see Appendix A in the Supplementary Material, along with, e.g., Kidziński et al., (2022), Wang and Carvalho, (2023), and Liu and Zhong, (2024). The choice of the parametrization typically depends on the specific application and the desired interpretation of the score and loading matrices. (B1) is the standard parametrization in principal component analysis, (B2) is the usual parametrization in spectral analysis of digital signals, and (B3) is the most common parametrization in the factor model literature. In our numerical experiments (Sections 5 and 6), we use parametrization (B1), which is conventional in the RNA-seq literature; see, e.g., Risso et al., (2018), Townes et al., (2019), and Ahlmann-Eltze and Huber, (2023).
@@ 2.2 Penalized likelihood estimation
$\sum_{i=1}^{n}\log\int_{\mathbb{R}^{d}}\Bigg[\prod_{j=1}^{m}f(y_{ij}\mid\mathbf{u}_{i:})\Bigg]f(\mathbf{u}_{i:})\,\textrm{{d}}\mathbf{u}_{i:},$ (4)
@@
with $\Psi$ being the parameter space of $\boldsymbol{\psi}$, which incorporate the identifiability constrains, and $\ell_{\lambda}(\boldsymbol{\psi};\mathbf{y})$ denoting the penalized negative log-likelihood function. The latter is given by
@@
where $\lambda>0$ is a regularization parameter and $\|\cdot\|_{F}^{2}$ denotes the Frobenius norm. Frobenius penalization is often introduced for numerical stability issues, but it is also intrinsically connected to the rank determination problem. Indeed, for any $\lambda>0$, the Frobenius penalty implicitly shrinks the singular values of the matrix factorization toward zero, encouraging compact low-rank representations of the signal. This spectral penalization effect and its connection with the nuclear norm have been extensively discussed in, e.g., Witten et al., (2009), Mazumder et al., (2010), and Kidziński et al., (2022).
@@ 2.3 Related models
Then, in the GMF notation, PCA can be obtained by assuming a Gaussian distribution for the data matrix, an identity link function for the mean and no regression effects in the linear predictor, namely $(\mathbf{X}\mathbf{B}^{\top})_{ij}=0$ and $(\boldsymbol{\Gamma}\mathbf{Z}^{\top})_{ij}=0$. Generalizations of PCA, such as Binary PCA (Schein et al.,, 2003; Lee et al.,, 2010; Landgraf,, 2015; Song et al.,, 2019) and Poisson PCA (Kenney et al.,, 2021; Virta and Artemiou,, 2023), are also included in the GMF framework.
@@ 3 Estimation algorithm
For the sake of exposition, throughout this section, we assume without loss of generality that $\mathbf{B}=\mathbf{0}$ and $\boldsymbol{\Gamma}=\mathbf{0}$. Moreover, we consider $\phi$ as a known fixed parameter, noting that, if unknown, it can be estimated iteration-by-iteration using either a method of moments or via maximum likelihood. More details about the general case $\mathbf{B}\neq\mathbf{0}$ and $\boldsymbol{\Gamma}\neq\mathbf{0}$ and the estimation of $\phi$ are provided in Appendix B in the Supplementary Material. Finally, we introduce the matrices $\accentset{\mbox{\large.}}{\mathbf{D}}=\{\accentset{\mbox{\large.}}{D}_{ij}\}=\{\partial\ell_{\lambda}/\partial\eta_{ij}\}$ and $\accentset{\mbox{\large.\hskip-1.33333pt.}}{\mathbf{D}}=\{\accentset{\mbox{\large.\hskip-1.33333pt.}}{D}_{ij}\}=\{\partial^{2}\ell_{\lambda}/\partial\eta_{ij}^{2}\}$ for the derivatives of the deviance with respect to the linear predictor, where
$\accentset{\mbox{\large.}}{D}(y,\mu)=\frac{\partial\ell_{\lambda}}{\partial\eta}=\frac{w(y-\mu)}{\phi\,\nu(\mu)\,\accentset{\mbox{\large.}}{g}(\mu)},\qquad\accentset{\mbox{\large.\hskip-1.33333pt.}}{D}(y,\mu)=\frac{\partial^{2}\ell_{\lambda}}{\partial\eta^{2}}=\frac{w\,\alpha(\mu)}{\phi\,\nu(\mu)\{\accentset{\mbox{\large.}}{g}(\mu)\}^{2}},$
and $\alpha(\mu)=1+(y-\mu)\{\accentset{\mbox{\large.}}{\nu}(\mu)/\nu(\mu)+\accentset{\mbox{\large.\hskip-1.33333pt.}}{g}(\mu)/\accentset{\mbox{\large.}}{g}(\mu)\}$. Accordingly, we define $\accentset{\mbox{\large.}}{g}=\textrm{{d}}{g}/\textrm{{d}}\mu$, $\accentset{\mbox{\large.\hskip-1.33333pt.}}{g}=\textrm{{d}}^{2}{g}/\textrm{{d}}\mu^{2}$ and $\accentset{\mbox{\large.}}{\nu}=\textrm{{d}}\nu/\textrm{{d}}\mu$. The expected second-order derivative, i.e., the Fisher weight $\mathbb{E}(\accentset{\mbox{\large.\hskip-1.33333pt.}}{D}_{ij})$, just corresponds to the observed second-order differential $\accentset{\mbox{\large.\hskip-1.33333pt.}}{D}_{ij}$ with $\alpha(\mu_{ij})=1$, which is positive for any $y_{ij}$ and $\mu_{ij}$.
@@
$\frac{\partial^{2}\ell_{\lambda}}{\partial\mathbf{U}^{2}}=\accentset{\mbox{\large.\hskip-1.33333pt.}}{\mathbf{D}}\,(\mathbf{V}*\mathbf{V})+\boldsymbol{\Lambda}>0,\qquad\frac{\partial^{2}\ell_{\lambda}}{\partial\mathbf{V}^{2}}=\accentset{\mbox{\large.\hskip-1.33333pt.}}{\mathbf{D}}^{\top}(\mathbf{U}*\mathbf{U})+\boldsymbol{\Lambda}>0,$ (8)
where * is the Hadamard product and $\boldsymbol{\Lambda}$ is a matrix of appropriate dimensions filled by $\lambda$.
Conditionally on $\mathbf{V}$, the left matrix equation in (7) can be decomposed into $n$ multivariate equations row-by-row, $\partial\ell_{\lambda}/\partial\mathbf{u}_{i:}=0$ ($i=1,\dots,n$), that can be solved independently in parallel. The existence and uniqueness of the solution of each row-equation are guaranteed under mild regularity conditions on the exponential family and the link function. In the same way, the right matrix equation in (7) can be split into $m$ independent vector equations, $\partial\ell_{\lambda}/\partial\mathbf{v}_{j:}=0$ ($j=1,\dots,m$), to be solved in parallel. See, e.g., Kidziński et al., (2022) and Wang and Carvalho, (2023) for a detailed discussion and derivation on (7) and (8). Notice that, in the presence of covariate effects, the derivatives in (7) must be replaced by $\partial\ell_{\lambda}/\partial[\boldsymbol{\Gamma},\,\mathbf{U}]$ and $\partial\ell_{\lambda}/\partial[\mathbf{B},\,\mathbf{V}]$, as detailed in Appendix B in the Supplementary Material.
@@ 3.1 Fisher scoring and quasi-Newton algorithms
The first, and most popular, algorithm introduced in the literature for finding the solution of (7) is the alternated iterative re-weighted least squares (AIRWLS, Collins et al.,, 2001; Li and Tao,, 2010; Risso et al.,, 2018; Kidziński et al.,, 2022; Wang and Carvalho,, 2023; Liu et al.,, 2023) method. It cycles between the conditional updates of $\mathbf{U}$ and $\mathbf{V}$ by solving the equations in (7) in a row-wise manner, using standard Fisher scoring for GLMs (McCullagh and Nelder,, 1989). The resulting routine is statistically motivated, easy to implement and allows for efficient parallel computing.
@@
where $\{\rho_{t}\}$ is a sequence of learning rate parameters, $\boldsymbol{\Delta}_{\psi}^{t}$ is the search direction, while $\mathbf{G}_{\psi}^{t}=\partial\ell_{\lambda}^{t}/\partial\boldsymbol{\psi}$ and $\mathbf{H}_{\psi}^{t}=\partial^{2}\ell_{\lambda}^{t}/\partial\boldsymbol{\psi}^{2}$ denote respectively the first two derivatives of $\ell_{\lambda}(\boldsymbol{\psi};\mathbf{y})$ with respect to $\boldsymbol{\psi}$ evaluated at $\boldsymbol{\psi}^{t}$. Throughout, $\leftarrow$ stands for the assignment operator and the division is performed element-wise. Exploiting the block structure of $\boldsymbol{\psi}$, the joint update (9) can be written in the coordinate-wise form
@@
where $\mathbf{G}_{\textsc{u}}$, $\mathbf{G}_{\textsc{v}}$, $\mathbf{H}_{\textsc{u}}$ and $\mathbf{H}_{\textsc{v}}$ can be obtained as in (7) and (8) (Kidziński et al.,, 2022). Overall, each quasi-Newton iteration requires $O((n+m)d+nmd)$ floating point operations and $O((n+m)d+nm)$ memory allocations. To the best of our knowledge, this is the most efficient algorithm in the literature for the estimation of GMF models.